How many milliliters of 0.260 M Na2S are needed to react with 40.00 mL of 0.315 M AgNO3, according to the following balanced equation? (7 points)

Na2S (aq) + 2 AgNO3 (aq) → 2 NaNO3 (aq) + Ag2S (s)
1
Expert's answer
2020-04-10T13:38:20-0400

According to the reaction we make the equation:

n(Na2S) = n(AgNO3)/2


n(AgNO3) = C( AgNO3)*V( AgNO3) = 0.315*0.04 = 0.0126 mol

n(Na2S)=0.0126/2=0.063 mol

V(Na2S)=n(Na2S)/C(Na2S)=0.063/0.26=0.2423 L=242.3 mL

Answer: 242.3 mL